# How Consistent Are Humans When Grading Programming Assignments?

### v1

@@ Abstract.
Our results show that human graders in our study can not agree on the grade to give a piece of student work and are often individually inconsistent, suggesting that the idea of a “gold standard” of human grading might be flawed, and highlights that a shared rubric alone is not enough to ensure consistency.
@@ 1. Introduction
Summative grades and formative feedback majorly impact the students’ outcomes throughout and after their degree. During their degree, their grades in the previous year’s courses can determine the students’ progression to the next year. Students’ final degree classification or grades are often factors in their employment prospects outside of education (Stepanova et al., 2021). Having processes to ensure consistent or fair marking is crucial to maintaining the perceived value of summative grades and students’ perceptions of justice(Nesbit and Burton, 2006).
While small cohorts are often assessed by a single assessor, such as the module/class leader, larger cohorts require multiple assessors, typically graduate teaching assistants, to provide grades and meaningful feedback in a short timeframe. These assessors often use a rubric comprising the evaluation criteria, quality definitions, and a scoring strategy (Reddy and Andrade, 2010) to grade and provide meaningful feedback on the students’ work. The rubrics can be written at varying granularities, from generic criteria encompassing the final grades, to individual rubric items allocating individual marks. However, using multiple assessors can lead to issues with how consistently the rubric is applied (Mustapha et al., 2016), especially for subjective elements of an assignment like code quality. Feedback quality can also vary between assessors, especially those who were themselves taught at different institutions or who have gained professional experience. Though rubrics provide a scoring strategy for teaching assistants, many experience uncertainty when assessing, even those who are experienced teaching assistants (Riese, 2018).
@@ 2.2. Teaching Assistants and Assessment
Teaching assistants are often crucial in supporting students’ learning; they typically consist of undergraduate, master’s and PhD students (Riese et al., 2021).
The three most common duties for teaching assistants include assisting in programming labs, leading sections to reinforce concepts, and grading; many institutions utilise undergraduate and graduate teaching assistants (Mirza et al., 2019).
Mirza et al. (2019) conducted a systematic literature review into how undergraduate teaching assistants (UTAs) are used within computer science courses and found that UTAs often grade assignments, whether they provide the final grades for the assignment, assign coarse primary grades or are one of multiple UTAs to grade each student assignment.
@@ 3.1.1. Assignment Outline
The assignment was a small-group, open-ended paired programming assignment to utilise object-oriented programming concepts to develop a predator/prey simulator with groups of two or three. The students were provided with a template project based on the “foxes-and-rabbit” project from the sixth edition of Barnes and Kölling (2006)’s book on “Objects First with Java”. The template includes a graphical user interface (GUI), a Field class, which contains a two-dimensional array for the simulation environment, and two animals, a Fox and a Rabbit; Figure 1 shows the class diagram of the template code. Students were asked to extend the template code with the following base tasks:
@@ 3.1.2. Dataset Limitations
The students were asked to conduct pair programming with explicit instructions on how to pair program. One student was the “driver”, who wrote the code, and the other was the “navigator”, who reviewed the code as it was written, and they were instructed to switch roles throughout the assignment. However, the assessment did not capture if and how the students conducted pair programming. Anecdotally, some groups started pair programming, especially in the associated lab sessions. However, as the submission deadline approached, groups tended to separate the work, with both students writing code in parallel.
@@ 3.2. Grading
To explore the consistency between graders, each assignment was graded four times, each time by a different grader. Figure 2 shows an example of what the data includes after the grading is complete. Grading each assignment multiple times by different graders allowed us to investigate how consistently a group of graders apply the grading rubric to individual assignments. To investigate self-consistency, we made graders grade a few assignments multiple times, spaced apart.
@@
In addition to collecting grades and feedback, we asked the graders to complete a diary each time they completed a grading session. The diary included where and when they graded, how long they graded, how many assignments they graded, and how rushed they felt to complete the grading that day. After completing the study, we invited the participants to complete a semi-structured interview, where we discussed their grading experience. Our prompt questions for the interview include:
@@
3.3. Analysis
To investigate how consistently a group of graders apply a grading rubric (RQ1), we used Krippendorff’s alpha (implemented using the Fast Krippendorff library (Castro, 2017)) to assess the reliability of agreement between the graders on the awarded grades. Krippendorff’s alpha is a generalised measure allowing any number of raters and categories while allowing for missing data (Krippendorff, 2013). It also has several weightings that can be applied to different types of data, including ordinal, which is data that is both categorical and has an order, such as letter grades. While the number of grades per assignment was fixed after participant recruitment, each assignment was only graded by four of the graders rather than all graders, making Krippendorff’s alpha the ideal measure. We evaluated the individual graders’ consistency using repeated assignments by validating the distance between the grades they awarded. This varies from our pre-registration, as our original choice of metric (Fliess’ Kappa) required a complete set of grades and did not consider the ordinal nature of grading.
To provide further context for how the graders undertook their grading, we analysed and produced summary statistics from the data provided by the grading diary, including when and where they conducted their grading, how long they graded, and how rushed they felt to complete the grading by the deadline. Additionally, we coded our semi-structured interview transcripts, which were automatically generated by Microsoft Teams and then validated by the authors.
@@ 4. Results
Our results section first explores the grader participant demographics, diary and interview results, which provides context for the analysis of the consistency of a group of graders before finally analysing the consistency of individual graders.
@@ 4.1. Participant Demographics
We conducted a pre-study survey to identify suitable grader participants and collect demographic data. We selected 28 participants to complete the grading.
Figure 3 shows the overall demographics of our participants. Most of our participants identified as male, with only five identifying as female, which follows a similar distribution of students enrolled in the computer science courses at our institution. We sought participants undertaking a PhD or Masters or who were in their final undergraduate year, with most participants being Masters students. Many of our participants studied or are studying their undergraduate at institutions in our country. While they may not have grading experience, they are well-versed in how grading is conducted at the university level in our country from the student perspective. These are a relatively typical profile of those recruited to act as graders at our institution, and as described in subsection 2.2, typical of graders at many other institutions as well.
(a) Gender
(b) Age
(c) Undergraduate Location
(d) Student Status
(e) Ethnicity
Figure 3. The demographics of our 28 participants, including gender, age, location of their undergraduate degree, student status and ethnicity.
Figure 4 shows how long our participants have been grading, if they have had previous grading experience, and how often they graded per academic year. Though most of our participants have never done grading before, all are experienced programmers, with only two participants programming for less than five years and all having some knowledge of Java. Figure 5 shows the time our participants have been programming and which languages they deem themselves proficient.
All participants received a participant expense payment congruent with hourly grading rates at our institution.
(a) Time Spent Grading
(b) Frequency of Grading
Figure 4. The grading experience of our participants, including how many years they have been grading and how often they graded per academic year.
(a) Time Programming
(b) Proficient Programming Languages
Figure 5. The programming experience of our participants, including how many years they have programmed and the programming languages they are proficient in.
@@
Figure 6 shows the degree level composition for each group of four. Nearly every group contained one or more PhD students, over half the groups contained two or more master’s students, and most groups and one or two third-year students.
4.2. Diary Results
In this section, we discuss the results of the grading diary, including how often the participants graded, how many assignments they graded per session, and details about the environment and time of day they graded.
Figure 12 shows a box plot of the number of diary entries per participant. Some participants opted to grade in two sessions, while others graded over 14, with the average grading over 6 sessions.
Figure 12 shows the relationship between the number of assignments graded in a session and the number of hours spent grading and shows that more time spent grading correlates to the number of assignments graded. However, a few graders do not follow this trend. Some graders take less time to grade many assignments, and one grader takes much longer to grade 6 assignments.
As part of the diary, we asked participants how rushed they felt to complete the grading, the results of which are in Figure 12. Most graders did not feel rushed to complete the grading, even with a two-week deadline to complete the grading and other external factors such as coursework deadlines and work.
Figures 12, 12, and 12 show the results of when, where and the environmental factors, respectively. Most grading sessions occurred at home in the evening while sitting silently or listening to music. While most grading was conducted in the evening, many sessions were conducted at other times of the day, with the number of sessions in morning (35) and night (34) nearly identical.
Figure 7. A box plot showing the summary of the number of diary entries per participant.
Figure 8. The count of assignments graded against the number of hours graded per session, as reported by the participants.
Figure 9. The results of how rushed the participants felt while grading. Some participants selected multiple values on the 5-point Likert scale, so we calculated the median for these values.
Figure 10. The count of the time of day when the participants reported they were grading.
Figure 11. The count of the locations where the participants reported they were grading.
Figure 12. The count of the environmental factors around the participants while they were grading.
4.3. Interview Results
We extracted key terms from our end-of-study semi-supervised interviews with our participants; Figure 13 shows the results. Most participants split each batch of twenty assignments into smaller groups (3 - 5 assignments) and graded each of the four elements before moving to the next assignment. Few opted to split the batch into larger groups (10+ assignments), and some graded each element before moving on to the next.
Nearly all of our participants thought they consistently applied the rubric throughout both batches, while some thought they became more consistent over time. While not specifically instructed to, many participants opted to review the grades they awarded at the end of the batch or their session.
As part of our study investigating the participants’ individual consistency, we duplicated a single assignment between the first and second batches. This duplicate assignment was assigned a different assignment number and selected based on the batch’s median grade. During our interview, we asked participants if they noticed any duplicated assignments. Excluding those submissions that submitted the template code, six participants (1, 2, 10, 11, 20, 25) reported noticing the duplicates; 22 participants did not see that they were regrading a repeated assignment. One who noticed the duplicate opted to revise the grades in both batches to F because they believed it was plagiarism.
Figure 13. The count of the answers to the post-study interview.

### v2

@@ Abstract.
Our results show that human graders in our study can not agree on the grade to give a piece of student work and are often individually inconsistent, suggesting that the idea of a “gold standard” of human grading might be flawed, and highlights that a shared rubric alone is not enough to ensure consistency, and other aspects such as assessor training and alternative grading practices should be explored to improve the consistency of human grading further when grading programming assignments.
@@ 1. Introduction
Summative assessment, where the student is evaluated on what they have learnt during the semester or chapter (Glazer, 2014), majorly impacts the students’ outcomes throughout and after their degree (Knight, 2002). During their degree, their grades in the previous year’s courses can determine the students’ progression to the next year. Students’ final degree classification or grades are often factors in their employment prospects outside of education (Stepanova et al., 2021; Knight, 2002). Having processes to ensure consistent or fair marking is crucial to maintaining the perceived value of summative grades and students’ perceptions of justice(Nesbit and Burton, 2006).
While small cohorts are often assessed by a single assessor, such as the module/class leader, larger cohorts require multiple assessors, typically graduate teaching assistants, to provide grades and meaningful feedback in a short timeframe. These assessors often use a rubric comprising the evaluation criteria, quality definitions, and a scoring strategy (Reddy and Andrade, 2010) to grade and provide meaningful feedback on the students’ work. The rubrics can be written at varying granularities, from generic criteria encompassing the final grades, to individual rubric items allocating individual marks. However, using multiple assessors can lead to issues with how consistently the rubric is applied (Mustapha et al., 2016), especially for subjective elements of an assignment like code quality. Though rubrics provide a scoring strategy for teaching assistants, many experience uncertainty when assessing, even those who are experienced teaching assistants (Riese, 2018).
@@ 2.1. Consistency in Grading
Willey and Gardner (2010), investigated methods to ensure marking consistency when using multiple tutors to grade a second-year engineering course in design fundamentals. Each tutor was asked to grade their session tutorial of 32 students, and to achieve a consistent standard of marking between different tutors, they used double-blind marking and remarking. As part of the grading process, the tutors were arranged into small groups to discuss and regrade the assignments collaboratively. The course coordinator also provided two benchmark assignments from previous years that the tutors subsequently graded using specified criteria, effectively reducing the variability in grading between different tutors. While the study found only small variations between markings from different tutors, students still complained of a perceived lack of consistency, specifically being exacerbated by inconsistencies in how different tutors delivered feedback.
@@
2.2. Rubric Design
Rubrics are often used to make assessment expectations transparent (Brookhart, 2018), and they describe desirable qualities and common pitfalls in student work against a gradation of quality (Andrade, 2005). A rubric has three essential features: the evaluation criteria, the definitions of quality and a scoring strategy (Reddy and Andrade, 2010).
A review by Reddy and Andrade (2010) discussed rubric use in higher education. They present 20 articles on rubrics and include topics on student and instructor perceptions, use and academic performance, and rubrics’ validity and reliability. They found that instructors can have contrasting perceptions of using rubrics, with some believing that rubrics provide an objective basis for evaluation, while others resist integrating rubrics into their assessment. The review also investigated how a well-designed scoring rubric should alleviate inconsistencies in the grading process, both when rubrics are used by a single grader or a group of grader. A well-designed rubric should minimise errors due to grader training, grader feedback, and clarity in the criteria description. They further discuss how there is ample evidence of disagreement between assessors when using rubrics and stress the importance of grader training.
Jonsson and Svingby (2007) reviewed the reliability, validity and educational consequences of using rubrics in assessment. They analysed 75 studies and found that rubrics can improve the reliability of assessment if they are analytic, topic-specific, and complemented with exemplars or rater training, especially with open-ended assessment, as it is not always possible to limit the format of the assessment to achieve high levels of reliability without sacrificing the validity.
Rinne (2024)’s recent study aimed to identify sources of inconsistency in assessing undergraduate dissertations from a primary school teacher education programme. They identify that rubrics are often short and open to interpretation, resulting in inconsistent assessment, and discuss sources of inconsistency when grading, which include examiners’ interpretations of quality, and found that the same grade could be awarded for different reasons, and the same judgement could result in different grades.
To increase the agreement in teachers’ grading, Anders Jönsson and Hartell (2021) conducted a study into whether analytic or holistic rubrics enabled consistent assessment. Analytic assessment involves assessing different aspects of student performance, such as mechanics, grammar, style and organisation. On the other hand, holistic assessment means making an overall assessment that considers all criteria simultaneously (Anders Jönsson and Hartell, 2021). Anders Jönsson and Hartell’s study investigated analytic and holistic grading in both an English as a foreign language (EFL) and a mathematics high-school course. They found that analytic grading is slightly preferable to holistic assessment regarding teacher agreement for both EFL and mathematics. They further conclude that the results support the argument that teachers’ grading is ‘complex, intuitive and tacit’ (Bloxham et al., 2016) and that any attempts to achieve grading consistency are likely futile. However, their results also indicate a small step towards increasing agreement.
In addition to the active area of research investigating how to design rubrics to maximise consistency while maintaining validity, within computer science, defining rubrics for aspects of programming, especially code quality, is still a developing field. Stegeman et al. (2016) defined a code quality rubric for introductory programming, including names, comments, formatting and expressions. They tested this rubric against several programming assignments and found that no substantial parts of code quality appeared to be missing. Recently, Kirk et al. (2024b), motivated by the difficulty for programming educators to find appropriate guidance on teaching style concepts, have defined a literature-informed model for code style, which includes many of the same topics introduced by Stegeman et al. but provides further depth and rationale behind the elements. The model provides an abstract level of the main style aspects that map to understanding and changing code and include principles that relate to language, layout, constructs, design, redundancy, implementation, repetition and structure.
@@ 2.3. Teaching Assistants and Assessment
Teaching assistants are often crucial in supporting students’ learning; they typically consist of undergraduate, master’s and PhD students (Riese et al., 2021; Wald and Harland, 2018; Riese and Bälter, 2022).
The three most common duties for teaching assistants include assisting in programming labs, leading sections to reinforce concepts, and grading; many institutions utilise undergraduate and graduate teaching assistants (Mirza et al., 2019; Willey and Gardner, 2010; Wald and Harland, 2018).
Mirza et al. (2019) conducted a systematic literature review into how undergraduate teaching assistants are used within computer science courses and found that UTAs often grade assignments, whether they provide the final grades for the assignment, assign coarse primary grades or are one of the multiple undergraduate teaching assistants to grade each student assignment.
@@
Wald and Harland (2018) discuss the trade-offs of employing student teaching assistants in their position paper, including how they can reduce teaching loads for academics and provide graduate students with teaching experience and financial support. However, teaching experience can differ between teaching assistants, especially undergraduate teaching assistants, who typically have limited academic and pedagogical knowledge, which are both essential for student learning. They suggest limiting the role of graduate and late-stage undergraduate teaching assistants as graders, tutors, and demonstrators, as they have yet to find any sound pedagogical arguments supporting undergraduate teaching assistants as graders. Instead, they suggest that they facilitate discussion and review course material.
Kristiansen et al. (2024) also highlight the importance of properly trained teaching assistants in their study comparing manual and automated feedback. They found in a controlled study of undergraduate students in a programming assignment that there was a need for both automated and manual feedback, providing superior results in task effectiveness and student preferences than just automated or manual assessment.
Riese and Bälter (2022) conducted a qualitative analysis into how course coordinators use assessment throughout their course and what they expect from TAs. They found that the seven experienced course coordinators in CS1 courses typically use TAs to provide continuous assessments for large cohorts. They discuss how some TAs set the bar too high while others set it too low, which can lead to consistency and that some TAs can view their students as friends, leading to ethical dilemmas. Similarly, to Kristiansen et al., Riese and Bälter suggest that not all tasks are suitable for TAs and that assessing students’ competency and knowledge can be difficult for experienced teachers.
@@ 3.1.1. Assignment Outline
The assignment was a small-group, open-ended paired programming assignment to utilise object-oriented programming concepts to develop a predator/prey simulator with groups of two or three. The students were asked to conduct pair programming with explicit instructions on how to pair program. One student was the “driver”, who wrote the code, and the other was the “navigator”, who reviewed the code as it was written, and they were instructed to switch roles throughout the assignment. The students were provided with a template project based on the “foxes-and-rabbit” project from the sixth edition of Barnes and Kölling (2006)’s book on “Objects First with Java”. The template includes a graphical user interface (GUI), a Field class, which contains a two-dimensional array for the simulation environment, and two animals, a Fox and a Rabbit; Figure 1 shows the class diagram of the template code. Students were asked to extend the template code with the following base tasks:
@@ 3.1.2. Dataset Limitations
While the students were given instructions on how to pair program, the assessment did not capture if and how the students conducted pair programming. Anecdotally, some groups started pair programming, especially in the associated lab sessions. However, as the submission deadline approached, groups tended to separate the work, with both students writing code in parallel.
@@ 3.2. Assessment
3.3. Rubric
The rubric we provided participants can be found in Appendix B. This is the original rubric from the assignment from where the submissions were captured. The participants were asked to supply letter grades between A and F, with ‘+’ and ‘-’ grades for all letters; for exceptional submissions, participants could award A++. Participants were asked to grade the following programming skills:
  • Correctness – Assess if a student has comprehended and completed the assignments in a way that complies with the assignment requirements (Gerdes et al., 2010).
  • Code Elegance – Investigate the overall code design. Have the students use Java design principles correctly, including inheritance and polymorphism, methods to reduce duplicated code, and their overall approach to implementing the required features (Börstler et al., 2018; Kirk et al., 2024b).
  • Readability – Evaluate whether a student’s submission is easy to understand. This can include code style, whether formatting, indentation and naming conventions (Börstler et al., 2018; Kirk et al., 2024b).
  • Documentation – Assess if and how well students have documented their code, including inline comments and the natural language code summary (docstring) (Börstler et al., 2018; Rai et al., 2022; Kirk et al., 2024b).
While code elegance, readability, and documentation are all aspects of code quality, we asked participants to provide grades for each aspect, allowing for richer data capture of different aspects of code quality. Similar aspects have been utilised in multiple different studies (Messer et al., 2024; Neutens et al., 2022)
This rubric consists of the three essential features of a rubric, as discussed in Section 2.2, including the evaluation criteria, the definitions of quality, and a scoring strategy. While some improvements could be made to the clarity of the definitions of quality and the scoring strategy, the rubric we provided our participants was the original rubric from the course, which provided the following:
[…]
  • The same rubric that historical students provided alongside the assignment requirements guided the design and development of their submissions.
4. Methodology
4.1. Group Consistency (RQ1)
To explore the consistency between graders, each assignment was graded four times, each time by a different grader. Figure 2 shows an example of what the data includes after the grading is complete. Grading each assignment multiple times by different graders allowed us to investigate how consistently a group of graders apply the grading rubric to individual assignments.
@@ 4.1. Group Consistency (RQ1)
To investigate how consistently a group of graders apply a grading rubric, we used Krippendorff’s $\alpha$ (implemented using the Fast Krippendorff library (Castro, 2017)) to assess the reliability of agreement between the graders on the awarded grades. Krippendorff’s $\alpha$ is a percent chance-based reliability coefficient, similar to Cohen’s $\kappa$ and Fleiss’ $\kappa$. It evaluates the deviation from perfect reliability by the proportion of observed to expected disagreement and allows any number of raters and categories while allowing for missing data (Krippendorff, 2013). It also has several weightings that can be applied to different types of data, including ordinal, which is data that is both categorical and has an order, such as letter grades.
As many percent chance agreement-based coefficients, including Krippendorf’s $\alpha$ and Fleiss’ $\kappa$, underestimate the agreement when one of the labels is dominant (Gwet, 2008; Doughty et al., 2024; Wongpakaran et al., 2013), we also compute Gwet’s $AC_{2}$, which is a weighted variant to Gwet’s $AC_{1}$, which was designed to be robust in the high prevalence of a single class (Gwet, 2008). Gwet’s $AC_{2}$ is more robust in instances of high prevalence, as it removes from consideration all agreements that occurred by chance (Gwet, 2008).
As we planned, each assignment was graded by four of the graders rather than all graders, making Krippendorff’s $\alpha$ and Gwet’s $AC_{2}$ the ideal measures. Using Krippendorf’s $\alpha$ and Gwet’s $AC_{2}$ varies from our pre-registration, as our original choice of metric (Fliess’ $\kappa$) required a complete set of grades and did not consider the ordinal nature of grading, so we realised that this was not the best choice.
4.2. Individual Consistency (RQ2)
@@
To evaluate the consistency of individual participants, we calculate the distance between the grade awarded in the first and second batches for each participant and for each skill. We could not use typical inter-rater analysis to evaluate self-consistency, as the two grades are too few for inter-rater reliability metrics, including Krippendorf’s $\alpha$, Gwet’s $AC_{2}$ and Fliess’ $\kappa$. We chose to limit the number of repeated assignments to ensure that the participants were unaware that they were grading a duplicate and to maximise the number of annotations for the dataset.
To provide further information on the participants’ experience grading these assignments, we conducted a semi-structured interview. These interviews were conducted in English by one of the members of the research team and lasted approximately 30 minutes; Microsoft Teams was used to provide automatic transcription. During the interview, the participants were asked the following questions:
@@
Unfortunately, many participants did not provide detailed answers for most of the questions asked, even when prompted for more information. This is likely due to the interviewer’s lack of experience conducting semi-structured interviews. As such, we have decided not to provide an in-depth analysis of the semi-structured interviews and only use the answers from questions four and nine in our results. These questions were coded as “yes” or “no” by one member of the research team and used to provide further context to our analysis of participant’s individual consistency.
@@ 4.3. Pilot Study
4.4. Participant Selection
We conducted a pre-study survey to identify suitable grader participants and collect demographic data. Our criteria for our participants follow the typical demographics of teaching assistants performing grading in many institutions (Mirza et al., 2019; Riese et al., 2021); and are as follows:
  • Students at our institutions in their third year of undergraduate, a master’s student or a PhD student.
  • They identify as proficient in Java programming and have at least three years of programming experience.
All selected participants were remunerated £280 GBP (\$340 USD) for participating; this rate was concurrent with our institution’s hourly rate for teaching assistants. We decided it was important that our participants receive pay similar to what they would receive when grading for a live assignment to ensure they take the study seriously and complete all their grading.
4.5. Position Statement
The researchers in this study are three academic staff members and a graduate student at a large UK university. The dataset is captured from one of the academic staff’s courses, CS1 Java courses that ran between 2017 and 2021 inclusive, with permission from the college’s ethics committee (LRS/DP-22/23-34272), the students’ permission was not required for this data, and all data was validated by research assistants to ensure the data was anonymised, as prescribed by the college’s data handling policy. The participants who took part in the study are current students at our institution but are not current students of any members of the research team. The consistency study was further approved by our college’s ethics committee to use the existing dataset and to provide remuneration to our participants (LRS/DP-22/23-35578).
@@ 5.1. Participant Demographics
Through our pre-study survey, we selected 28 participants who matched our criteria to complete the study. 25 of our participants had over five years of programming experience, with three having between three and five years, and all identified as proficient in Java. Seven participants were PhD students, 13 were Masters students, and eight were third-year undergraduate students. Two participants graded for between two and three years, one participant graded between one and two years, six participants graded for less than 1 year, and 21 participants had not been assessed before. These are a relatively typical profile of those recruited to act as graders at our institution, and as described in subsection 2.3, typical of graders at many other institutions as well.
Most of our participants identified as male, with three identifying as female, which follows a similar distribution of students enrolled in the computer science courses at our institution.
Figure 3 shows the degree level composition for each group of four. Nearly every group contained one or more PhD students, over half the groups contained two or more master’s students, and most groups had one or two third-year students.